### v1

# The $\Lambda_{b}\to\Lambda$ transition form factors in perturbative QCD approach

Abstract
In this work, we investigate the $\Lambda_{b}\to\Lambda$ transition form factors in the perturbative QCD (PQCD) approach, incorporating higher-twist light-cone distribution amplitudes (LCDAs). The resulted form factors show that higher-twist LCDAs are dominant numerically. By combining our PQCD predictions at low-$q^{2}$ with lattice QCD results at high-$q^{2}$, $z$-series expansion fits are performed to parametrize the form factors over the full kinematic range. We also provide the prediction for physical observables in the rare decay $\Lambda_{b}\to\Lambda\mu^{+}\mu^{-}$, including the differential branching fraction, dilepton longitudinal polarization fraction, and forward-backward asymmetries (lepton-side, hadron-side, and combined lepton-hadron). Our obtained form factors are consistent with those in other theoretical methods within the uncertainties.
I introduction
In the standard model (SM), flavor-changing neutral currents (FCNCs) such as the $b\to s$ transition are strongly suppressed by the Glashow-Iliopoulos-Maiani (GIM) mechanism Glashow:1970gm , as they can only proceed through loop diagrams. However, new physics scenarios could potentially enhance these FCNC processes by several orders of magnitude. Making FCNC decays excellent probes for testing the SM and searching for new physics. The LHCb Collaboration has observed several deviations between theoretical expectations and experimental measurements in $b\to sl^{+}l^{-}$ transitions, including $B^{+}\to K^{(*)+}\mu^{+}\mu^{-}$, $B^{0}\to K^{0}\mu^{+}\mu^{-}$ LHCb:2014cxe , $B^{0}\to K^{*0}\mu^{+}\mu^{-}$ LHCb:2016ykl , $B^{0}_{s}\to\phi\mu^{+}\mu^{-}$ LHCb:2021zwz , and the baryonic decay $\Lambda_{b}^{0}\to\Lambda\mu^{+}\mu^{-}$ LHCb:2015tgy . These observed anomalies motivate our theoretical investigation on the transition form factors.
While rare semileptonic decays of B mesons have been extensively studied, the corresponding baryonic decays remain less explored. This work focuses specifically on the $\Lambda_{b}\to\Lambda$ transition, aiming to provide complementary information to B meson decays. The heart of the $\Lambda_{b}\to\Lambda$ transition lies the accurate determination of the form factors, which play a pivotal role in theoretical predictions, such as the differential branching fractions of semileptonic decays and the CP-related observables in non-leptonic processes. Various theoretical approaches have been employed to investigate the $\Lambda_{b}\to\Lambda$ form factors, including soft-collinear effective theory Feldmann:2011xf , lattice QCD calculations Detmold:2012vy ; Detmold:2016pkz , and QCD light-cone sum rules Wang:2015ndk , each offering distinct advantages in addressing the challenging problem.
Among various theoretical approaches, the perturbative QCD (PQCD) approach, preserving the transverse momentum based on the $k_{T}$ factorization, has proven effective for studying heavy-to-light hadronic transitions in the fast recoil region Kurimoto:2001zj ; Li:2012nk . This approach is especially valuable as it can evaluate nonfactorizable contributions unambiguously, such as the $W$-exchange and penguin-annihilation amplitudes, which play a crucial role in predicting physical observables for nonleptonic decays, particularly the CP violation (CPV) Lu:2000em ; Cheng:2014rfa . Thus, the PQCD approach has been successfully employed to predict the CP violations of the $B^{0}\to K^{+}\pi^{-}$ and $B^{0}\to\pi^{+}\pi^{-}$ decays Belle:2004mad ; BaBar:2004gyj ; Belle:2004nch , and recently explain the long-standing puzzle of why the measured direct CP violations in $\Lambda_{b}\to p\pi^{-},pK^{-}$ are so small compared to their $B$ meson counterparts LHCb:2018fly ; LHCb:2024iis ; Han:2024kgz ; Han:2025tvc .
Recent measurements have significantly advanced our understanding of CP violation in $\Lambda_{b}$ baryon decays. The LHCb collaboration firstly observed baryon CP violation in the four-body decay $\Lambda_{b}\to pK^{-}\pi^{+}\pi^{-}$, and reported a large local CP asymmetry exceeding $6.0\sigma$ significance in the region of $m_{p\pi^{+}\pi^{-}}<2.7\text{GeV}/c^{2}$ LHCb:2025ray . It implies the importance of dynamical studies of $\Lambda_{b}$ decays by the consistency between the experimental result in LHCb:2025ray and the theoretical prediction within the $N\pi$ rescattering mechanism Wang:2024oyi . Besides, the interference between resonances is instrumental in generating CP violation for multibody decays Wang:2024oyi ; Yu:2025ekh . The dominant decay modes of $\Lambda$ baryon to nucleon and pion provide substantial strong phases that may enhance local CP violation in cascade processes. The LHCb Collaboration also found evidence of CP violation in the three-body decay $\Lambda_{b}\to\Lambda K^{+}K^{-}$ LHCb:2024yzj . These experimental breakthroughs motivate our systematic investigation of the $\Lambda_{b}\to\Lambda$ transition form factors within the PQCD framework, aiming to establish a robust theoretical foundation for interpreting these intriguing observations in the future works.
The first application of the PQCD approach to form factor calculations was carried out in Refs. Li:1992ce ; Kundu:1998gv , where the proton electromagnetic form factors were computed, incorporating Sudakov suppression. Subsequent studies in Refs. Shih:1998pb ; Lu:2009cm extended this approach to the hadronic transition $\Lambda_{b}\to p$, finding that the resulting form factors were significantly smaller than those obtained from nonperturbative methods. Recently, Ref. Han:2022srw confirmed the validity of the PQCD approach by including subleading-twist baryon LCDAs, successfully reproducing the $\Lambda_{b}\to p$ form factors with a reasonable magnitude.
The form factors of the VAV-A current for the $\Lambda_{b}\to\Lambda$ transition were studied in Ref. Rui:2022sdc , where different parameter sets for the $\Lambda$ baryon LCDAs were examined. The parameters derived from lattice QCD Bali:2015ykx ; RQCD:2019hps only account for leading-twist LCDAs, leading to excessively small form factors. Another choice is from the light-cone QCD sum rules Liu:2014uha , which incorporates high-power corrections up to twist-6, including next-to-leading-order conformal spin expansion contributions. However, the resulting form factors are anomalously large and violate heavy-quark symmetry, suggesting that the $\Lambda$ baryon LCDA calculations given in Ref. Liu:2014uha may contain problematic assumptions.
In this work, we systematically investigate all $10$ form factors describing the $\Lambda_{b}\to\Lambda$ transition induced by the $b\to s$ vector, axial-vector, and tensor currents in the PQCD approach. The $\Lambda$ baryon LCDAs with high-power corrections up to twist-6 but with leading order conformal spin expansion accuracy Liu:2009mb are employed to incorporate subleading power contributions. Our calculations successfully reproduce reasonable $\Lambda_{b}\to\Lambda$ form factors, in agreement with those from other theoretical approaches. We also demonstrate the subleading power contributions dominate baryon transition form factors, which is consistent with the conclusion in Ref. Han:2022srw .
The remainder of this paper is organized as follows. Sect. II summarizes the definitions of the $\Lambda_{b}\to\Lambda$ transition form factors and presents the PQCD framework for baryon transition calculations. The LCDAs for the $\Lambda_{b}$ and $\Lambda$ baryons used in our analysis are detailed in Sect. III. We then discuss our numerical results for the form factors and their extrapolation using the modified z-expansion in Sect. IV. Sect. V presents our predictions for the differential branching fraction and angular observables in $\Lambda_{b}\to\Lambda(\to p\pi^{-})\mu^{+}\mu^{-}$ decays. Finally, we conclude with a summary in Sect. VI. The Appendix provides the explicit factorization formulas for each relevant diagram.
II Definition Of The Form Factors and PQCD framework
The helicity-based definition of the $\Lambda_{b}\to\Lambda$ form factors are given by Feldmann:2011xf
$\langle\Lambda(p^{\prime},s^{\prime})|\overline{s}\gamma^{\mu}b|\Lambda_{b}(p,s)\rangle=$ $\overline{u}_{\Lambda}(p^{\prime},s^{\prime})\Big{[}f_{0}(q^{2})\,(m_{\Lambda_{b}}-m_{\Lambda})\frac{q^{\mu}}{q^{2}}$
$+f_{+}(q^{2})\frac{m_{\Lambda_{b}}+m_{\Lambda}}{s_{+}}\left(p^{\mu}+p^{\prime\mu}-(m_{\Lambda_{b}}^{2}-m_{\Lambda}^{2})\frac{q^{\mu}}{q^{2}}\right)$
$+f_{\perp}(q^{2})\left(\gamma^{\mu}-\frac{2m_{\Lambda}}{s_{+}}p^{\mu}-\frac{2m_{\Lambda_{b}}}{s_{+}}p^{\prime\mu}\right)\Big{]}u_{\Lambda_{b}}(p,s)$ (1)
$\langle\Lambda(p^{\prime},s^{\prime})|\overline{s}\gamma^{\mu}\gamma_{5}b|\Lambda_{b}(p,s)\rangle=$ $-\overline{u}_{\Lambda}(p^{\prime},s^{\prime})\gamma_{5}\Big{[}g_{0}(q^{2})\,(m_{\Lambda_{b}}+m_{\Lambda})\frac{q^{\mu}}{q^{2}}$
$+g_{+}(q^{2})\frac{m_{\Lambda_{b}}-m_{\Lambda}}{s_{-}}\left(p^{\mu}+p^{\prime\mu}-(m_{\Lambda_{b}}^{2}-m_{\Lambda}^{2})\frac{q^{\mu}}{q^{2}}\right)$
$+g_{\perp}(q^{2})\left(\gamma^{\mu}+\frac{2m_{\Lambda}}{s_{-}}p^{\mu}-\frac{2m_{\Lambda_{b}}}{s_{-}}p^{\prime\mu}\right)\Big{]}u_{\Lambda_{b}}(p,s)$ (2)
$\langle\Lambda(p^{\prime},s^{\prime})|\overline{s}\,i\sigma^{\mu\nu}q_{\nu}\,b|\Lambda_{b}(p,s)\rangle=$ $-\overline{u}_{\Lambda}(p^{\prime},s^{\prime})\Big{[}h_{+}(q^{2})\frac{q^{2}}{s_{+}}\left(p^{\mu}+p^{\prime\mu}-(m_{\Lambda_{b}}^{2}-m_{\Lambda}^{2})\frac{q^{\mu}}{q^{2}}\right)$
$+h_{\perp}(q^{2})(m_{\Lambda_{b}}+m_{\Lambda})\left(\gamma^{\mu}-\frac{2m_{\Lambda}}{s_{+}}p^{\mu}-\frac{2m_{\Lambda_{b}}}{s_{+}}p^{\prime\mu}\right)\Big{]}u_{\Lambda_{b}}(p,s)$ (3)
$\langle\Lambda(p^{\prime},s^{\prime})|\overline{s}\,i\sigma^{\mu\nu}q_{\nu}\gamma_{5}\,b|\Lambda_{b}(p,s)\rangle=$ $-\overline{u}_{\Lambda}(p^{\prime},s^{\prime})\gamma_{5}\Big{[}\widetilde{h}_{+}(q^{2})\frac{q^{2}}{s_{-}}\left(p^{\mu}+p^{\prime\mu}-(m_{\Lambda_{b}}^{2}-m_{\Lambda}^{2})\frac{q^{\mu}}{q^{2}}\right)$
$+\widetilde{h}_{\perp}(q^{2})(m_{\Lambda_{b}}-m_{\Lambda})\left(\gamma^{\mu}+\frac{2m_{\Lambda}}{s_{-}}p^{\mu}-\frac{2m_{\Lambda_{b}}}{s_{-}}p^{\prime\mu}\right)\Big{]}u_{\Lambda_{b}}(p,s)$ (4)
with $p$ being the momentum of the $\Lambda_{b}$ baryon and $p^{\prime}$ the momentum of the $\Lambda$ baryon. $q=p-p^{\prime}$ denotes the momentum transferred. $\sigma^{\mu\nu}=\frac{i}{2}(\gamma^{\mu}\gamma^{\nu}-\gamma^{\nu}\gamma^{\mu})$ and $s_{\pm}=(m_{\Lambda_{b}}\pm m_{\Lambda})^{2}-q^{2}$. In some other analysis, the above matrix elements are usually decomposed into the first and second class form factors according to the Weinberg classification El-Khadra:2001wco ; Gutsche:2013pp :
$\langle\Lambda(p^{\prime},s^{\prime})|\overline{s}\,\gamma^{\mu}\,b|\Lambda_{b}(p,s)\rangle=$ $\overline{u}_{\Lambda}(p^{\prime},s^{\prime})\left[f_{1}^{V}(q^{2})\>\gamma^{\mu}-\frac{f_{2}^{V}(q^{2})}{m_{\Lambda_{b}}}i\sigma^{\mu\nu}q_{\nu}+\frac{f_{3}^{V}(q^{2})}{m_{\Lambda_{b}}}q^{\mu}\right]u_{\Lambda_{b}}(p,s),$ (5)
$\langle\Lambda(p^{\prime},s^{\prime})|\overline{s}\,\gamma^{\mu}\gamma_{5}\,b|\Lambda_{b}(p,s)\rangle=$ $\overline{u}_{\Lambda}(p^{\prime},s^{\prime})\left[f_{1}^{A}(q^{2})\>\gamma^{\mu}-\frac{f_{2}^{A}(q^{2})}{m_{\Lambda_{b}}}i\sigma^{\mu\nu}q_{\nu}+\frac{f_{3}^{A}(q^{2})}{m_{\Lambda_{b}}}q^{\mu}\right]\gamma_{5}u_{\Lambda_{b}}(p,s),$ (6)
[…]
The relation between these two set form factors are given by Detmold:2016pkz ,
$f_{+}(q^{2})$ $=$ $f_{1}^{V}(q^{2})+\frac{q^{2}}{m_{\Lambda_{b}}(m_{\Lambda_{b}}+m_{\Lambda})}f_{2}^{V}(q^{2}),$ (9)
$f_{\perp}(q^{2})$ $=$ $f_{1}^{V}(q^{2})+\frac{m_{\Lambda_{b}}+m_{\Lambda}}{m_{\Lambda_{b}}}f_{2}^{V}(q^{2}),$ (10)
$f_{0}(q^{2})$ $=$ $f_{1}^{V}(q^{2})+\frac{q^{2}}{m_{\Lambda_{b}}(m_{\Lambda_{b}}-m_{\Lambda})}f_{3}^{V}(q^{2}),$ (11)
$g_{+}(q^{2})$ $=$ $f_{1}^{A}(q^{2})-\frac{q^{2}}{m_{\Lambda_{b}}(m_{\Lambda_{b}}-m_{\Lambda})}f_{2}^{A}(q^{2}),$ (12)
$g_{\perp}(q^{2})$ $=$ $f_{1}^{A}(q^{2})-\frac{m_{\Lambda_{b}}-m_{\Lambda}}{m_{\Lambda_{b}}}f_{2}^{A}(q^{2}),$ (13)
$g_{0}(q^{2})$ $=$ $f_{1}^{A}(q^{2})-\frac{q^{2}}{m_{\Lambda_{b}}(m_{\Lambda_{b}}+m_{\Lambda})}f_{3}^{A}(q^{2}),$ (14)
$h_{+}(q^{2})$ $=$ $-f_{2}^{TV}(q^{2})-\frac{m_{\Lambda_{b}}+m_{\Lambda}}{m_{\Lambda_{b}}}f_{1}^{TV}(q^{2}),$ (15)
$h_{\perp}(q^{2})$ $=$ $-f_{2}^{TV}(q^{2})-\frac{q^{2}}{m_{\Lambda_{b}}(m_{\Lambda_{b}}+m_{\Lambda})}f_{1}^{TV}(q^{2}),$ (16)
$\widetilde{h}_{+}(q^{2})$ $=$ $-f_{2}^{TA}(q^{2})+\frac{m_{\Lambda_{b}}-m_{\Lambda}}{m_{\Lambda_{b}}}f_{1}^{TA}(q^{2}),$ (17)
$\widetilde{h}_{\perp}(q^{2})$ $=$ $-f_{2}^{TA}(q^{2})+\frac{q^{2}}{m_{\Lambda_{b}}(m_{\Lambda_{b}}-m_{\Lambda})}f_{1}^{TA}(q^{2}).$ (18)
In the framework of PQCD approach, the transition matrix elements of the $\Lambda_{b}\to\Lambda$ can be evaluated as the convolution of the hard kernels and baryon LCDAs,
$\mathcal{A}\sim\Psi_{\Lambda_{b}}\left(x_{i},b_{i},\mu\right)\otimes\mathcal{H}\left(x_{i},b_{i},x^{\prime}_{i},b^{\prime}_{i},t,\mu\right)\otimes\Psi_{\Lambda}\left(x^{\prime}_{i},b^{\prime}_{i},\mu\right),$ (19)
where $\Psi_{\Lambda_{b}}$ and $\Psi_{\Lambda}$ are the LCDAs of the $\Lambda_{b}$ baryon and $\Lambda$ baryon, respectively, which are non-perturbative and serve as the main sources of error in our calculation. The hard kernel $\mathcal{H}$ is derived perturbatively from diagrams. $x_{i}$ and $x^{{}^{\prime}}_{i}$ are the momentum fractions carried by quarks of $\Lambda_{b}$ and $\Lambda$, $b_{i}\left(b^{{}^{\prime}}_{i}\right)$ are the conjugate variables of the quarks’ transverse momentum $k_{iT}\left(k^{{}^{\prime}}_{iT}\right)$, respectively. The hard scale $t$ denotes the maximal virtual scale of the internal propagator within the hard kernel. Refer to Li:2001ay ; Lu:2000hj ; Han:2022srw for more details of the PQCD framework and its application in baryon decays.
In the rest frame of the $\Lambda_{b}$ baryon, the momentum of the $\Lambda_{b}$ baryon and $\Lambda$ baryon are parameterized in the light-cone coordinate as,
$p=\frac{m_{\Lambda_{b}}}{\sqrt{2}}(1,1,\vec{\textbf{0}}),\quad p^{\prime}=\frac{m_{\Lambda_{b}}}{\sqrt{2}}(\eta_{1},\eta_{2},\vec{\textbf{0}}),$ (20)
with
$\eta_{1}=\frac{M^{2}_{\Lambda_{b}}+m^{2}_{\Lambda}-q^{2}\pm\sqrt{\left(M^{2}_{\Lambda_{b}}+m^{2}_{\Lambda}-q^{2}\right)^{2}-4M^{2}_{\Lambda_{b}}m^{2}_{\Lambda}}}{2M^{2}_{\Lambda_{b}}},$
$\eta_{2}=\frac{M^{2}_{\Lambda_{b}}+m^{2}_{\Lambda}-q^{2}\mp\sqrt{\left(M^{2}_{\Lambda_{b}}+m^{2}_{\Lambda}-q^{2}\right)^{2}-4M^{2}_{\Lambda_{b}}m^{2}_{\Lambda}}}{2M^{2}_{\Lambda_{b}}},$ (21)
where $M_{\Lambda_{b}}$ and $m_{\Lambda}$ denote the masses of the $\Lambda_{b}$ and $\Lambda$ baryons, respectively. The partonic momentum with the transverse momentum retained for the quarks are parameterized as,
$k_{1}$ $=\left(\frac{m_{\Lambda_{b}}}{\sqrt{2}},\frac{x_{1}m_{\Lambda_{b}}}{\sqrt{2}},\vec{\textbf{k}}_{1T}\right),$ $k_{1}^{\prime}$ $=\left(x_{1}^{\prime}\frac{\eta_{1}m_{\Lambda_{b}}}{\sqrt{2}},0,\vec{\textbf{k}}_{1T}^{\prime}\right),$
$k_{2}$ $=\left(0,\frac{x_{2}m_{\Lambda_{b}}}{\sqrt{2}},\vec{\textbf{k}}_{2T}\right),$ $k_{2}^{\prime}$ $=\left(x_{2}^{\prime}\frac{\eta_{1}m_{\Lambda_{b}}}{\sqrt{2}},0,\vec{\textbf{k}}_{2T}^{\prime}\right),$
$k_{3}$ $=\left(0,\frac{x_{3}m_{\Lambda_{b}}}{\sqrt{2}},\vec{\textbf{k}}_{3T}\right),$ $k_{3}^{\prime}$ $=\left(x_{3}^{\prime}\frac{\eta_{1}m_{\Lambda_{b}}}{\sqrt{2}},0,\vec{\textbf{k}}_{3T}^{\prime}\right),$ (22)
$k_{1}$ is the momentum of $b$ quark, $k_{1}^{{}^{\prime}}$ is the momentum of $s$ quark, and $k_{2}$ ($k_{3}$), $k_{2}^{\prime}$ ($k_{3}^{\prime}$) are the spectator $u$ ($d$) quark’s momentum in the $\Lambda_{b}$ and the $\Lambda$ baryons, respectively. $x_{i}^{(\prime)}$ and $k_{T}^{(\prime)}$ denote the corresponding momentum fractions and the the transverse momentum. In the $\Lambda_{b}$ baryon, $b$ quark carries most of the momentum fraction, $x_{1}\sim{\cal O}\left(m_{b}^{2}/M^{2}_{\Lambda_{b}}\right)$, where $m_{b}$ is the mass of b quark. $u,d$ quarks carry small momentum fractions, being the soft quarks that require at least two hard gluons to transmit momentum to them in order to form energetic quarks in the $\Lambda$ baryon, as shown in Fig. 1. In other words, ${\cal O}\left(\alpha_{s}^{2}\right)$ is the leading order for calculating the transition amplitude of $\Lambda_{b}\to\Lambda$ in the PQCD approach.
Figure 1: The leading-order Feynman diagrams for the $\Lambda_{b}\to\Lambda$ transition form factors in the PQCD approach. The green $\otimes$ denote the weak interaction currents. These diagrams are labelled by $d_{1},d_{2},\dots,d_{16}$.
III Light-cone distribution amplitudes
As stated before, the necessary inputs in the PQCD approach to calculate the $\Lambda_{b}\to\Lambda$ transition form factors are the baryon LCDAs. The LCDAs of $\Lambda_{b}$ baryon are defined by the hadron matrix elements of the non-local operators sandwiched between the vacuum state and the $\Lambda_{b}$ baryon state. The general Lorentz structures in the momentum-space are Wang:2015ndk ; Bell:2013tfa ; Ball:2008fw
$\begin{split}(Y_{\Lambda_{b}})_{\alpha\beta\gamma}(x_{i},\mu)&\equiv\frac{1}{2\sqrt{2}N_{c}}\int\prod_{l=2}^{3}\frac{dz_{l}^{-}d\bm{z}_{l}}{(2\pi)^{3}}e^{ik_{l}\cdot z_{l}}\epsilon^{ijk}\langle 0|T[b_{\alpha}^{i}(0)u_{\beta}^{j}(z_{2})d_{\gamma}^{k}(z_{3})]|\Lambda_{b}\rangle\\ &=\frac{1}{8\sqrt{2}N_{c}}\Big{\{}f_{\Lambda_{b}}^{(1)}(\mu)[M_{1}(x_{2},x_{3})\gamma_{5}C^{T}]_{\gamma\beta}+f_{\Lambda_{b}}^{(2)}(\mu)[M_{2}(x_{2},x_{3})\gamma_{5}C^{T}]_{\gamma\beta}\Big{\}}[\Lambda_{b}(p)]_{\alpha},\end{split}$ (23)
where $\alpha,\beta,\gamma$ are the spinor indices, $i,j,k$ are the color indices and $N_{c}$ is the number of colors. $C^{T}$ represents the transpose of the charge conjugate matrix, and $\Lambda_{b}\left(p\right)$ represrnts the spinor of the $\Lambda_{b}$ baryon. $f_{\Lambda_{b}}^{\left(1\right)}$ and $f_{\Lambda_{b}}^{\left(2\right)}$ are normalized constants, here we choose $f_{\Lambda_{b}}^{\left(1\right)}\approx f_{\Lambda_{b}}^{\left(2\right)}\equiv f_{\Lambda_{b}}=0.03\pm 0.005GeV^{3}$ Han:2022srw ; Rui:2022sdc . The remaining parts of the chiral-even projector $M_{1}$ and the chiral-odd projector $M_{2}$ are expressed as,
[…]
where $n$ and $\bar{n}$ are two light-cone vectors, $n=\left(1,0,\vec{\textbf{0}}\right)$, $\bar{n}=\left(0,1,\vec{\textbf{0}}\right)$, satisfying $n\cdot\bar{n}=1$. Several theoretical models for the $\Lambda_{b}$ baryon LCDA functions $\Psi_{2},\Psi_{3}^{+-},\Psi_{3}^{-+},\Psi_{4}$ have been proposed in Refs. Ball:2008fw ; Bell:2013tfa ; Ali:2012zza . In this work we adopt the exponential mode from Bell:2013tfa which is in a simple form and considered to be good enough in current stage Han:2022srw ; Rui:2022sdc . The exponential mode specific as follows Bell:2013tfa
$\Psi_{2}(x_{2},x_{3})$ $=$ $x_{2}x_{3}\frac{M_{\Lambda_{b}}^{4}}{\omega_{0}^{4}}e^{-\frac{(x_{2}+x_{3})M_{\Lambda_{b}}}{\omega_{0}}},$
$\Psi_{3}^{+-}(x_{2},x_{3})$ $=$ $2x_{2}\frac{M_{\Lambda_{b}}^{3}}{\omega_{0}^{3}}e^{-\frac{(x_{2}+x_{3})M_{\Lambda_{b}}}{\omega_{0}}},$
$\Psi_{3}^{-+}(x_{2},x_{3})$ $=$ $2x_{3}\frac{M_{\Lambda_{b}}^{3}}{\omega_{0}^{3}}e^{-\frac{(x_{2}+x_{3})M_{\Lambda_{b}}}{\omega_{0}}},$
$\Psi_{4}(x_{2},x_{3})$ $=$ $\frac{M_{\Lambda_{b}}^{2}}{\omega_{0}^{2}}e^{-\frac{(x_{2}+x_{3})M_{\Lambda_{b}}}{\omega_{0}}},$ (25)
where the parameter $\omega_{0}$ denotes the average energy of the two light quarks and is chosen to be $0.7\pm 0.1$GeV in the PQCD approach as suggested by the calculation of the $\Lambda_{b}\to p\pi^{-},pK^{-}$ decays Han:2024kgz ; Han:2025tvc .
In the SU(3) flavor symmetry limit, the LCDAs of the final-state $\Lambda$ baryon up to twist-6 have been defined through the corresponding momentum-space projector in Liu:2009mb ; Liu:2014uha ,
[…]
The light-like vector $\mathcal{P^{\prime}}$ is defined, by the $\Lambda$ baryon momentum $p^{\prime}$ and the light-like vector $z$, as
$\mathcal{P^{\prime}}_{\mu}=p^{\prime}_{\mu}-\frac{1}{2}z_{\mu}\frac{m_{\Lambda}^{2}}{\mathcal{P^{\prime}}\cdot z},$ (27)
[…]
  • twist-3
    $V_{1}(x_{i})$ $=$ $0\,,\hskip 128.0374ptA_{1}(x_{i})=-120x_{1}x_{2}x_{3}\phi_{3}^{0}\,$,
    $T_{1}(x_{i})$ $=$ $0\,.$ (28)
  • twist-4
    $S_{1}(x_{i})$ $=$ $6x_{3}(1-x_{3})(\xi_{4}^{0}+\xi_{4}^{{}^{\prime}0})\,,\hskip 42.67912ptP_{1}(x_{i})=6x_{3}(1-x_{3})(\xi_{4}^{0}-\xi_{4}^{{}^{\prime}0})\,$,
    $V_{2}(x_{i})$ $=$ $0\,,\hskip 142.26378ptA_{2}(x_{i})=-24x_{1}x_{2}\phi_{4}^{0}\,$,
    $V_{3}(x_{i})$ $=$ $12(x_{1}-x_{2})x_{3}\psi_{4}^{0}\,,\hskip 65.44142ptA_{3}(x_{i})=-12x_{3}(1-x_{3})\psi_{4}^{0}\,$,
    $T_{2}(x_{i})$ $=$ $0\,,\hskip 142.26378ptT_{3}(x_{i})=6(x_{2}-x_{1})x_{3}(-\xi_{4}^{0}+\xi_{4}^{{}^{\prime}0})\,$,
    $T_{7}(x_{i})$ $=$ $-6(x_{1}-x_{2})x_{3}(\xi_{4}^{0}+\xi_{4}^{{}^{\prime}0})\,.$ (29)
  • twist-5
    $S_{2}(x_{i})$ $=$ $\frac{3}{2}(x_{1}+x_{2})(\xi_{5}^{0}+\xi_{5}^{{}^{\prime}0})\,,\hskip 42.67912ptP_{2}(x_{i})=\frac{3}{2}(x_{1}+x_{2})(\xi_{5}^{0}-\xi_{5}^{{}^{\prime}0})\,$,
    $V_{4}(x_{i})$ $=$ $3(x_{2}-x_{1})\psi_{5}^{0}\,,\hskip 79.6678ptA_{4}(x_{i})=-3(1-x_{3})\psi_{5}^{0}\,$,
    $V_{5}(x_{i})$ $=$ $0\,,\hskip 139.41832ptA_{5}(x_{i})=-6x_{3}\phi_{5}^{0}\,$,
    $T_{4}(x_{i})$ $=$ $-\frac{3}{2}(x_{1}-x_{2})(\xi_{5}^{0}+\xi_{5}^{{}^{\prime}0})\,,\hskip 36.98866ptT_{5}(x_{i})=0\,$,
    $T_{8}(x_{i})$ $=$ $-\frac{3}{2}(x_{1}-x_{2})(\xi_{5}^{0}-\xi_{5}^{{}^{\prime}0})\,.$ (30)
  • twist-6
    $V_{6}(x_{i})$ $=$ $0\,,\hskip 128.0374ptA_{6}(x_{i})=-2\phi_{6}^{0}\,$,
    $T_{6}(x_{i})$ $=$ $0\,.$ (31)
The values $\phi_{i}^{0},\psi_{i}^{0},\xi_{i}^{0}$ and $\xi_{i}^{\prime 0}$ in Eqs. (III-III) connect to four independent parameters as follows,
$\phi_{3}^{0}$ $=$ $\phi_{6}^{0}=-f_{\Lambda},\hskip 79.6678pt\phi_{4}^{0}=\phi_{5}^{0}=-\frac{1}{2}(f_{\Lambda}+\lambda_{1}),$
$\psi_{4}^{0}$ $=$ $\psi_{5}^{0}=\frac{1}{2}(f_{\Lambda}-\lambda_{1}),\hskip 45.52458pt\xi_{4}^{0}=\xi_{5}^{0}=\lambda_{2}+\lambda_{3},$
$\xi_{4}^{{}^{\prime}0}$ $=$ $\xi_{5}^{{}^{\prime}0}=\lambda_{3}-\lambda_{2}.$ (32)
Their numerical values and uncertainties have been derived in QCD sum rules Liu:2009mb ,
$f_{\Lambda}$ $=$ $(6.0\pm 0.3)\times 10^{-3}\;\mbox{GeV}^{2},\hskip 71.13188pt\lambda_{1}=(1.0\pm 0.3)\times 10^{-2}\;\mbox{GeV}^{2},$
$|\lambda_{2}|$ $=$ $(0.83\pm 0.05)\times 10^{-2}\;\mbox{GeV}^{2},\hskip 56.9055pt|\lambda_{3}|=(0.83\pm 0.05)\times 10^{-2}\;\mbox{GeV}^{2}.$ (33)
It is noting that the parameters $\lambda_{2}$ and $\lambda_{3}$ in the above parameters are only specifies their absolute values. In the following calculation, we adopt four scheme depending on different choices of the signs of these two parameters.
  • Scheme-I: $\lambda_{2}=(0.83\pm 0.05)\times 10^{-2}\mbox{GeV}^{2}$,  $\lambda_{3}=(0.83\pm 0.05)\times 10^{-2}\mbox{GeV}^{2}$.
  • Scheme-II: $\lambda_{2}=(0.83\pm 0.05)\times 10^{-2}\mbox{GeV}^{2}$,  $\lambda_{3}=(-0.83\pm 0.05)\times 10^{-2}\mbox{GeV}^{2}$.
  • Scheme-III: $\lambda_{2}=(-0.83\pm 0.05)\times 10^{-2}\mbox{GeV}^{2}$,  $\lambda_{3}=(0.83\pm 0.05)\times 10^{-2}\mbox{GeV}^{2}$.
  • Scheme-IV: $\lambda_{2}=(-0.83\pm 0.05)\times 10^{-2}\mbox{GeV}^{2}$,  $\lambda_{3}=(-0.83\pm 0.05)\times 10^{-2}\mbox{GeV}^{2}$.
IV numerical results
The PQCD factorization formula for each feynman diagrams are collected in Appendix. Appendix A:Factorization formulas. We present the numerical results of the form factors at $q^{2}=0\>\text{GeV}^{2}$ under four different schemes in Table 1, comparing them with lattice QCD calculations Detmold:2016pkz . The uncertainties in our calculated results originate from the shape parameter in the exponential model LCDAs of $\Lambda_{b}$ baryon and parameters in the $\Lambda$ baryon LCDAs, The more accurate parameters in the baryon LCDAs will help us to further improve the precision of our theoretical predictions. The results in Table 1 show that Scheme-I and Scheme-II are in good agreement with the lattice QCD predictions using the nominal form factor parameters. In particular, Scheme-I also exhibits excellent consistency with the lattice results obtained using the higher-order parameterization. In contrast, Scheme-III and Scheme-IV yield systematically smaller form factors. This implies that the parameter $\lambda_{2}$ plays an important role in correctly reproducing the magnitude of $\Lambda_{b}\to\Lambda$ form factors, whereas $\lambda_{3}$ has less impact. For simplicity, we therefore adopt Scheme-I in the following calculations and analyses.
Table 1: Form factors of $\Lambda_{b}\to\Lambda$ obtained via four schemes, compared with lattice QCD results Detmold:2016pkz , where [1] and [2] represent lattice results with nominal and higher-order parameter choices, respectively.
Scheme-I Scheme-II Scheme-III Scheme-IV lattice QCD[1] Detmold:2016pkz lattice QCD[2] Detmold:2016pkz
$f_{+}(0)$ $0.192\pm 0.058$ $0.130\pm 0.043$ $114\pm 0.040 0.0$$049\pm 0.030$ $0.160\pm 0.024$ $0.202\pm 0.040$
$f_{\perp}(0)$ $0.239\pm 0.072$ $0.161\pm 0.053$ $142\pm 0.050 0.0$$061\pm 0.036$ $0.207\pm 0.037$ $0.241\pm 0.065$
$f_{0}(0)$ $0.192\pm 0.058$ $0.129\pm 0.043$ $114\pm 0.040 0.0$$049\pm 0.029$ $0.156\pm 0.035$ $0.200\pm 0.041$
$g_{+}(0)$ $0.194\pm 0.059$ $0.132\pm 0.044$ $117\pm 0.041 0.0$$053\pm 0.030$ $0.112\pm 0.028$ $0.197\pm 0.036$
$g_{\perp}(0)$ $0.165\pm 0.050$ $0.113\pm 0.038$ $100\pm 0.036 0.0$$46\pm 0.026 0.0$$095\pm 0.035$ $0.169\pm 0.055$
$g_{0}(0)$ $0.195\pm 0.060$ $0.132\pm 0.044$ $117\pm 0.041 0.0$$053\pm 0.030$ $0.166\pm 0.023$ $0.199\pm 0.038$
$h_{+}(0)$ $0.239\pm 0.072$ $0.161\pm 0.053$ $142\pm 0.050 0.0$$061\pm 0.035$ $0.236\pm 0.041$ $0.228\pm 0.072$
$h_{\perp}(0)$ $0.193\pm 0.059$ $0.131\pm 0.044$ $115\pm 0.041 0.0$$051\pm 0.030$ $0.166\pm 0.022$ $0.219\pm 0.043$
$\widetilde{h}_{+}(0)$ $0.165\pm 0.051$ $113\pm 0.038 0.0$$99\pm 0.036 0.0$$045\pm 0.027$ $0.163\pm 0.029$ $0.263\pm 0.050$
$\widetilde{h}_{\perp}(0)$ $0.193\pm 0.059$ $0.131\pm 0.044$ $116\pm 0.041 0.0$$051\pm 0.030$ $0.156\pm 0.027$ $0.235\pm 0.044$
Table 1 presents our calculated helicity form factors, which satisfy the expected endpoint relations $f_{0}(0)=f_{+}(0)$ and $g_{0}(0)=g_{+}(0)$. For comparison with other theoretical approaches, we convert these to Weinberg form factors in Table 2. These results are consistent with the heavy quark effective theory predictions, where the Weinberg form factors for heavy-to-light baryon transitions obey:
$f_{1}^{V}(0)$ $\approx f_{1}^{A}(0)\gg f_{2,3}^{V,A}(0),$
$f_{2}^{TV}(0)$ $\approx f_{2}^{TA}(0)\gg f_{1}^{TV,TA}(0).$ (34)
as clearly demonstrated in Table 2. Within theoretical uncertainties, our form factor calculations agree well with results from other approaches.
Table 2: Theoretical predictions for the form factors at $q^{2}=0$ of the $\Lambda_{b}\rightarrow\Lambda$ transition, where the symbol $\cdots$ denotes that no relevant result has been reported in the literature.
$f_{1}^{V}(0)$ $f_{2}^{V}(0)$ $f_{3}^{V}(0)$ $f_{1}^{A}(0)$ $f_{2}^{A}(0)$ $f_{3}^{A}(0)$
this work $192\pm 0.058 0.0$$39\pm 0.011 0.0$$035\pm 0.010$ $195\pm 0.060 0.0$$37\pm 0.011 0.0$$037\pm 0.011$
Lattice QCD[1]Detmold:2016pkz $0.158$ $0.041$ 0.017-0.017 $0.137$ $0.055$ 0.117-0.117
Lattice QCD[2]Detmold:2016pkz $0.201$ $0.033$ 0.039-0.039 $0.198$ $0.036$ 0.166-0.166
LCSRWang:2015ndk $0.18$ $0.017$ $\cdots$ $0.18$ 0.025-0.025 $\cdots$
LCSRAliev:2010uy $322\pm 0.112 0.0$$62\pm 0.022 0.0$$084\pm 0.028$ $318\pm 0.110 0.0$$73\pm 0.022 0.0$$079\pm 0.028$
relativistic QMFaustov:2017wbh $0.208$ $0.302$ $0.026$ $0.125$ 0.003-0.003 0.083-0.083
covariant QMGutsche:2013oea $0.107$ $0.043$ $0.003$ $0.104$ $0.003$ 0.052-0.052
QCDSRHuang:1998ek $0.446$ $0.073$ 0.073-0.073 $0.446$ $0.073$ 0.073-0.073
PQCDRui:2022sdc $0.095^{+0.057+0.018}_{-0.029-0.021}$ $\cdots$ $\cdots$ $0.104^{+0.060+0.016}_{-0.033-0.020}$ $\cdots$ $\cdots$
QCDFZhu:2018jet $131_{-0.018-0.008}^{+0.016+0.008} 0.0$$48_{-0.008-0.002}^{+0.008+0.002} 0.0$$027_{-0.006-0.000}^{+0.006+0.000}$ $132_{-0.017-0.009}^{+0.016+0.008} 0.0$$24_{-0.005-0.001}^{+0.005+0.001} 0.0$$052_{-0.008-0.001}^{+0.008+0.001}$
NRQMMott:2011cx $0.025$ $\cdots$ $\cdots$ $0.028$ $\cdots$ $\cdots$
LFQMWei:2009np $0.108$ $0.031$ 0.031-0.031 $0.106$ $0.006$ 0.006-0.006
GFAMohanta:2000nk $0.061$ $0.025$ 0.008-0.008 $0.107$ $0.014$ 0.043-0.043
NRQMCheng:1996cs $0.062$ $0.025$ 0.008-0.008 $0.108$ $0.014$ 0.043-0.043
$f^{TV}_{1}(0)$ $f^{TV}_{2}(0)$ $f^{TA}_{1}(0)$ $f^{TA}_{2}(0)$
this work $0.038\pm 0.011$ $193\pm 0.059 0.0$$036\pm 0.010$ $0.193\pm 0.059$
Lattice QCD[1]Detmold:2016pkz 0.059-0.059 0.166-0.166 $0.009$ 0.156-0.156
Lattice QCD[2]Detmold:2016pkz 0.008-0.008 0.219-0.219 $0.034$ 0.235-0.235
LCSRWang:2015ndk 0.025-0.025 0.18-0.18 $0.037$ 0.18-0.18
LCSRWang:2015ndk 0.025-0.025 0.18-0.18 $0.037$ 0.18-0.18
LCSRAliev:2010uy $0.070\pm 0.022$ $295\pm 0.105 0.0$$091\pm 0.032$ $0.294\pm 0.105$
relativistic QMFaustov:2017wbh 0.029-0.029 0.153-0.153 $0.029$ 0.153-0.153
covariant QMGutsche:2013oea 0.043-0.043 0.105-0.105 $0.003$ 0.105-0.105
QCDSRHuang:1998ek 0.076-0.076 0.446-0.446 0.083-0.083 0.446-0.446
Table 3 presents the numerical results of the form factors $f_{1}^{V,A}(0)$ decomposed by individual Feynman diagrams. Our analysis reveals that diagrams $d_{7}$, $d_{8}$, $d_{13}$, and $d_{14}$ provide the dominant contributions to $f_{1}^{V,A}(0)$.
Table 3: All diagram results for the form factors $f_{1}^{V}$ and $f_{1}^{A}$ at $q^{2}=0$ of the $\Lambda_{b}\rightarrow\Lambda$ transition.
$d_{1}$ $d_{2}$ $d_{3}$ $d_{4}$ $d_{5}$ $d_{6}$ $d_{7}$ $d_{8}$
$f_{1}^{V}(0)$ $1.7\times 10^{-3}$ $1.7\times 10^{-3}$ $2.6\times 10^{-3}$ $1.1\times 10^{-4}$ $9.5\times 10^{-4}$ $7.2\times 10^{-4}$ 0.051 0.044
$f_{1}^{A}(0)$ $1.8\times 10^{-3}$ $1.8\times 10^{-3}$ $2.8\times 10^{-3}$ $1.6\times 10^{-4}$ $1.2\times 10^{-3}$ $1.7\times 10^{-3}$ 0.051 0.045
$d_{9}$ $d_{10}$ $d_{11}$ $d_{12}$ $d_{13}$ $d_{14}$ total
$f_{1}^{V}(0)$ $1.1\times 10^{-3}$ $9.6\times 10^{-4}$ $5.3\times 10^{-3}$ $1.7\times 10^{-3}$ 0.047 0.053 0.192
$f_{1}^{A}(0)$ $1.1\times 10^{-3}$ $1.0\times 10^{-3}$ $5.4\times 10^{-3}$ $3.3\times 10^{-4}$ 0.048 0.055 0.195
We analyze the contributions of each twist-term combination for these four diagrams in Table 4. The diagrams $d_{7}$ and $d_{8}$ receive their primary contributions from the combination of $\Lambda_{b}$ twist-4 and $\Lambda$ twist-5 terms, and the diagrams $d_{13}$ and $d_{14}$ are predominantly governed by $\Lambda_{b}$ twist-3 and $\Lambda$ twist-4 combinations. These findings demonstrate the essential role of high-power effects in baryon transition form factors at the scale of the b-quark mass.
Table 4: The contribution of the diagrams $d_{7},d_{8},d_{13}$ and $d_{14}$ to the form factor $f_{1}^{V}(0)$.
$d_{7}$ $d_{8}$
$\Lambda$ twist-3 $\Lambda$ twist-4 $\Lambda$ twist-5 $\Lambda$ twist-6 $\Lambda$ twist-3 $\Lambda$ twist-4 $\Lambda$ twist-5 $\Lambda$ twist-6
$\Lambda_{b}$ twist-2 $\sim 0$ $\sim 0$ $\sim 0$ $\sim 0$ $\sim 0$ $\sim 0$ $\sim 0$ $\sim 0$
$\Lambda_{b}$ twist-3 $\sim 0$ $\sim 0$ $\sim 0$ $\sim 0$ $\sim 0$ $\sim 0$ $\sim 0$ $\sim 0$
$\Lambda_{b}$ twist-4 0.0078 $\sim 0$ 0.0427 $\sim 0$ 0.0690 $\sim 0$ 0.0372 $\sim 0$
$d_{13}$ $d_{14}$
$\Lambda_{b}$ twist-2 $\sim 0$ $\sim 0$ $\sim 0$ $\sim 0$ $\sim 0$ $\sim 0$ $\sim 0$ $\sim 0$
$\Lambda_{b}$ twist-3 $\sim 0$ 0.0308 $\sim 0$ $\sim 0$ $\sim 0$ 0.0351 $\sim 0$ $\sim 0$
$\Lambda_{b}$ twist-4 $\sim 0$ -0.0006 0.0164 $\sim 0$ $\sim 0$ -0.0007 0.0185 $\sim 0$
Recalling the form factors are expressed as the convolution of the perturbatively calculated hard kernels and the non-perturbative baryon LCDAs. We list the numerators of the hard kernels for the diagram $d_{7}$ contributing to the form factor $f_{1}^{V}(0)$ in Table 5, classified by different twist combination. $i=2,3^{+-},3^{-+},4$ stands for the $\Lambda_{b}$ baryon LCDA and $j=3,4,5,6$ denotes the $\Lambda$ baryon LCDA. The mass ratio, $r=m_{\Lambda}/M_{\Lambda_{b}}$, reflects the $1/m_{b}$ power suppression associated with the high-twist $\Lambda$ baryon LCDAs. By comparing the $h_{43}$ and $h_{45}$ terms in Table 5, it is found that $h_{45}$ is power suppressed by factor of $r^{2}$ compared to $h_{43}$. However, $h_{43}$ includes an additional momentum fraction $x_{3}$.
Fig. 2 displays the dependence of the $\Lambda_{b}$ baryon LCDAs on the momentum fractions $x_{1}$, $x_{2}$ and $x_{3}$ in the exponential model. The $\Lambda_{b}$ twist-4 LCDA $\Psi_{4}$ shows a pronounced peak when $x_{1}\approx 1$ and $x_{2,3}$ approach zero, clearly indicating that the b-quark carries most of the baryon’s momentum. The presence of a small momentum fraction $x_{3}$ closing to zero significantly suppresses the contribution of $h_{43}$. Similarly, near-vanishing momentum fraction $x_{3}$ also significantly suppresses the numerical results of $h_{3^{+-}4}$ and $h_{3^{-+}4}$. These results demonstrate that high-power contributions to the baryon transition form factors receive substantial endpoint enhancement, emerging as the dominant effect, which is in agreement with the conclusions of Ref. Wang:2011uv ; Han:2022srw .
Table 5: Numerators $h_{ij}$ of the integrands in the factorization formula for the form factor $f_{1}^{V}(0)$ with $i=2,3^{+-},3^{-+},4$ and $j=3,4,5,6$. an overall coefficient $\mathcal{C}m_{\Lambda_{b}}^{3}f_{\Lambda_{b}}/(128N_{c}^{2})$ with the color factor $\mathcal{C}=8/3$ is implicit for the entries.
$h_{ij}$ $\Lambda$ twist-3 $\Lambda$ twist-4 $\Lambda$ twist-5 $\Lambda$ twist-6
$\Lambda_{b}$ twist-2 $\sim 0$ $\sim 0$ $8r^{2}x_{3}\Psi_{2}\left(A_{4}+A_{5}+V_{5}-V_{4}\right)$ $\sim 0$
$\Lambda_{b}$ twist-$3^{+-}$ $\sim 0$ $rx_{3}\Psi_{3}^{+-}\left(P_{1}-S_{1}-T_{3}-T_{7}\right)$ $\sim 0$ $\sim 0$
$\Lambda_{b}$ twist-$3^{-+}$ $\sim 0$ $4rx_{3}\Psi_{3}^{-+}\left(P_{1}+S_{1}-T_{3}+T_{7}-2T_{2}\right)$ $\sim 0$ $\sim 0$
$\Lambda_{b}$ twist-4 $8x_{3}\Psi_{4}\left(A_{1}-V_{1}\right)$ $\sim 0$ $8r^{2}\left(1-x_{2}^{\prime}\right)\Psi_{4}\left(V_{4}-A_{4}-2A_{5}\right)$ $\sim 0$
Figure 2: Dependencies of the $\Lambda_{b}$ baryon LCDAs $\psi_{2}$, $\psi_{3}^{+-}$, $\psi_{3}^{-+}$ and $\psi_{4}$ on the momentum fractions $x_{i}$ for the exponential model in Bell:2013tfa . Each point inside the triangles satisfies the relation $x_{1}+x_{2}+x_{3}=1$.
To calculate the observables of semileptonic decays, we require the behavior of the $\Lambda_{b}\to\Lambda$ transition form factors across the whole kinematic range $0\leq q^{2}\leq q_{max}^{2}=(m_{\Lambda_{b}}-m_{\Lambda})^{2}$. The form factors are first evaluated at three $q^{2}$ points in the low-$q^{2}$ region, $0\leq q^{2}\leq m_{\tau}^{2}$, where PQCD approach is reliable, and then extended to the high-$q^{2}$ regime using the z-series formula Bourrely:2008za .
$F(q^{2})=\frac{1}{1-q^{2}/(m_{\rm pole}^{F})^{2}}\sum_{i=0}^{k}\left(a_{i}^{F}[z(q^{2})]^{i}\right).$ (35)
where $F\equiv f_{+,\perp,0},g_{+,\perp,0},h_{+,\perp},\widetilde{h}_{+,\perp}$, the parameter $z$ is defined as,
$z(q^{2})=\frac{\sqrt{t_{+}-q^{2}}-\sqrt{t_{+}-t_{0}}}{\sqrt{t_{+}-q^{2}}+\sqrt{t_{+}-t_{0}}},$ (36)
with $t_{0}=q^{2}_{\rm max}=(M_{\Lambda_{b}}-m_{\Lambda})^{2}$ and $t_{+}=(m_{B}+m_{K})^{2}$. The pole masses are taken from Detmold:2016pkz and collected in Table 6. We truncate the $z$-series expansion formula in Eq. (35) at $k=1$ for simplicity,
$F(q^{2})=\frac{1}{1-q^{2}/(m_{\rm pole}^{F})^{2}}\big{[}a_{0}^{F}+a_{1}^{F}\>z(q^{2})\big{]}.$ (37)
The central values and uncertainties of the parameters $a_{0}^{F}$ and $a_{1}^{F}$, obtained through a global fit using the three points at low-$q^{2}$ region calculated by the PQCD approach and one point at $q_{max}^{2}$ from the lattice QCD Detmold:2016pkz , are listed in Table 7. The form factors over the full kinematic region are presented in Fig. 3.
$f$ $J^{P}$ $m_{\rm pole}^{f}$ [GeV]
$_{+}, ff_{\perp}$, $_{+}, hh_{\perp}$ $1^{-}$ $5.416$
$f_{0}$ $0^{+}$ $5.711$
$_{+}, gg_{\perp}$, $\widetilde{h}_{+}$, $\widetilde{h}_{\perp}$ $1^{+}$ $5.750$
$g_{0}$ $0^{-}$ $5.367$
Table 6: Values of the pole masses, $m_{\rm pole}^{f}$ Detmold:2016pkz .
Parameter Value Parameter Value
$a_{0}^{f_{+}}$ $\phantom{-}0.419\pm 0.031$ $a_{0}^{g_{\perp}}$ $0.358\pm 0.027$
$a_{1}^{f_{+}}$ $1.04\pm 0.142$ $a_{1}^{g_{\perp}}$ $0.890\pm 0.123$
$a_{0}^{f_{0}}$ $\phantom{-}0.371\pm 0.032$ $a_{0}^{h_{+}}$ $\phantom{-}0.490\pm 0.050$
$a_{1}^{f_{0}}$ $0.830\pm 0.146$ $a_{1}^{h_{+}}$ $1.150\pm 0.208$
$a_{0}^{f_{\perp}}$ $\phantom{-}0.520\pm 0.040$ $a_{0}^{h_{\perp}}$ $\phantom{-}0.385\pm 0.034$
$a_{1}^{f_{\perp}}$ $1.290\pm 0.167$ $a_{1}^{h_{\perp}}$ $0.890\pm 0.156$
$a_{0}^{g_{+}}$ $\phantom{-}0.356\pm 0.023$ $a_{0}^{\widetilde{h}_{+}}$ $\phantom{-}0.339\pm 0.029$
$a_{1}^{g_{+}}$ $0.760\pm 0.107$ $a_{1}^{\widetilde{h}_{+}}$ $0.810\pm 0.134$
$a_{0}^{g_{0}}$ $\phantom{-}0.397\pm 0.029$ $a_{0}^{\widetilde{h}_{\perp}}$ $0.343\pm 0.029$
$a_{1}^{g_{0}}$ $\phantom{-}-0.930\pm 0.137$ $a_{1}^{\widetilde{h}_{\perp}}$ $0.700\pm 0.130$
Table 7: Central values and uncertainties of the nominal form factor parameters.
Figure 3: Form factors of the entire kinematic range, and the bands show the uncertainty.
V observables
The differential decay rate of the $\Lambda_{b}$ decay four-body $\Lambda_{b}\to\Lambda(\to p^{+}\pi^{-})\,\ell^{+}\ell^{-}$ with the $\Lambda_{b}$ unpolarized is derived as Gutsche:2013pp ; Boer:2014kda
$\frac{\mathrm{d}^{4}\Gamma}{\mathrm{d}q^{2}\>\mathrm{d}\cos\theta_{\ell}\>\mathrm{d}\cos\theta_{\Lambda}\>\mathrm{d}\phi}$ $=$ $\frac{3}{8\pi}\bigg{[}\hskip 12.05553pt\big{(}K_{1ss}\sin^{2}\theta_{\ell}+\,K_{1cc}\cos^{2}\theta_{\ell}+K_{1c}\cos\theta_{\ell}\big{)}$ (38)
[…]
where the angles $\theta_{\Lambda}$ and $\theta_{\ell}$ describe the polarization directions of the proton and negatively charged lepton, respectively. $K_{n\lambda}$ with $n=1,\dots,4$ and $\lambda=s,ss,c,cc,sc$ are functions of $q^{2}$, and its explicit expressions can be found in Eqs.(3.29)-(3.32) of Ref. Boer:2014kda . $\phi$ is the azimuthal angle between the decay planes of $p\pi$ and $\ell\ell$. By integrating over the three angles appearing in Eq. (38), we can obtain the $q^{2}$-differential decay rate as
$\frac{\mathrm{d}\Gamma}{\mathrm{d}q^{2}}=2K_{1ss}+K_{1cc}.$ (39)
In addition to the differential decay rate, we also define the normalized angular observables.
$\hat{K}_{i}=\frac{K_{i}}{\mathrm{d}\Gamma/\mathrm{d}q^{2}},$ (40)
Then the fraction of longitudinally polarized dileptons, the lepton-side forward-backward asymmetry, the hadron-side forward-backward asymmetry, and a combined lepton-hadron forward-backward asymmetry can be defined as
$F_{L}$ $=$ $2\hat{K}_{1ss}-\hat{K}_{1cc},$ (41)
$A_{\rm FB}^{\ell}$ $=$ $\frac{3}{2}\hat{K}_{1c},$ (42)
$A_{\rm FB}^{\Lambda}$ $=$ $\hat{K}_{2ss}+\frac{1}{2}\hat{K}_{2cc},$ (43)
$A_{\rm FB}^{\ell\Lambda}$ $=$ $\frac{3}{4}\hat{K}_{2c}.$ (44)
The parameters $\hat{K}_{2ss}$, $\hat{K}_{2cc}$, and $\hat{K}_{2c}$ in Eqs. (43) and (44) are related to the asymmetry parameters of $\Lambda\to p\pi^{-}$. Here, we adopt the values provided by the PDG ParticleDataGroup:2024cfk , $\alpha_{-}=0.747\pm 0.009$.
Figures 4 and 5 present our theoretical predictions for, respectively, the differential branching fraction of the $\Lambda_{b}\to\Lambda\mu^{+}\mu^{-}$ decay and the angular observables in the quasi four-body decay $\Lambda_{b}\to\Lambda(\to p\pi^{-})\mu^{+}\mu^{-}$. In the current analysis, we focus exclusively on the dominant theoretical uncertainty arising from the form factors, while other potential sources of uncertainty are neglected as they are expected to be subdominant in comparison. By integrating the differential branching fraction over the kinematic range $(2m_{\mu})^{2}\leq q^{2}\leq(M_{\Lambda_{b}}-m_{\Lambda})^{2}$, we obtain the total branching fraction for the $\Lambda_{b}\to\Lambda\mu^{+}\mu^{-}$ decay:
$\int_{(2m_{\mu})^{2}}^{(M_{\Lambda_{b}}-m_{\Lambda})^{2}}\frac{\mathrm{d}\mathcal{B}}{\mathrm{d}q^{2}}\mathrm{d}q^{2}=\left(1.24^{+0.25}_{-0.22}\right)\times 10^{-6}.$ (45)
where the integration is performed numerically across the full physical phase space. This prediction agrees well with the current experimental measurement of $(1.08\pm 0.28)\times 10^{-6}$ reported by the PDG ParticleDataGroup:2024cfk .
As evident from Fig. 4, the theoretical uncertainty in our result stemming from the form factor calculations is large. Future improvements in determining the light-cone distribution amplitudes (LCDAs) of both the $\Lambda_{b}$ and $\Lambda$ baryons could significantly reduce this uncertainty. The angular observables presented in Fig. 5, being defined as ratios of decay rates, exhibit less sensitivity to these uncertainties due to partial cancellations between numerator and denominator.
Figure 4: $\Lambda_{b}\to\Lambda\>\mu^{+}\mu^{-}$ differential branching fraction, compared to the experimental data from LHCb LHCb:2015tgy .
Figure 5: $\Lambda_{b}\to\Lambda(\to p^{+}\pi^{-})\mu^{+}\mu^{-}$ angular observables, compared to the experimental data from LHCb LHCb:2015tgy .
VI Summary
In this paper, we calculate the $\Lambda_{b}\to\Lambda$ transition form factors by the PQCD approach, and our results for the transition form factors in $q^{2}$ region are close to those in other theoretical approaches within errors. In our calculation, we consider the contributions of $\Lambda_{b}$ LCDAs up to twist-4 and $\Lambda$ LCDAs up to twist-6, We find that the combination term of $\Lambda_{b}$ twist-4 and $\Lambda$ twist-5 provides the largest contribution to the form factors, this indicates that the contributions of higher-twist terms in the LCDAs of baryons should be properly taken into account in the study of $\Lambda_{b}$ baryon decays. In addition, we obtained the form factors over the full kinematic region by performing a global fit that combines the PQCD results in the low-$q^{2}$ region with the lattice QCD results in the high-$q^{2}$ region. Using these form factors, we calculated the differential branching fraction and angular observables. Our results are consistent with most theoretical predictions and experimental data within the uncertainty range. Furthermore, due to the adoption of the latest $\Lambda_{b}\to p\pi$ asymmetry parameter, our prediction for hadron-side asymmetry is in better agreement with experimental data.
Acknowledgment
The authors would like to thank Dr. Dong-Hao Li for helpful discussions and valuable suggestions. The work is supported by the National Natural Science Foundation of China (Grant No. 12275067 and 12335003), Science and Technology R&D Program Joint Fund Project of Henan Province (Grant No. 225200810030), Science and Technology Innovation Leading Talent Support Program of Henan Province, National Key R&D Program of China (Grant No. 2023YFA1606000), and the Fundamental Research Funds for the Central Universities (Grant No. lzujbky-2023-stlt01, lzujbky-2024-oy02 and lzujbky-2025-eyt01).
Appendix A:Factorization formulas
We list below the factorization formulas for the form factor $f_{1}(q^{2}=0)$ from the Feynman diagrams $D_{i}$, i=1i=1-14, Those for the other five form factors can be derived in a similar way. The last two diagrams with three-gluon vertices do not contribute due to the vanishing color factors:
$f_{1}^{D_{1}}(q^{2}=0)=$ $\mathcal{C}M_{\Lambda_{b}}^{3}\frac{f_{\Lambda_{b}}}{8N_{c}}\frac{1}{8\sqrt{2}N_{c}}\int[dx]\int[dx^{\prime}]16\pi^{2}\alpha_{s}^{2}(t^{D_{1}})h_{f_{1}}\frac{1}{(2\pi)^{5}}\int b_{1}^{\prime}db_{1}^{\prime}\int b_{2}db_{2}$
$\int b_{3}db_{3}\int d\theta_{1}\int d\theta_{2}\exp[-S^{D_{1}}(x,x^{\prime},b,b^{\prime})]F_{1}(D,b_{3})F_{3}(A,B,C,b_{1}^{\prime},b_{2}),$ (46)
$f_{1}^{D_{2}}(q^{2}=0)=$ $\mathcal{C}M_{\Lambda_{b}}^{3}\frac{f_{\Lambda_{b}}}{8N_{c}}\frac{1}{8\sqrt{2}N_{c}}\int[dx]\int[dx^{\prime}]16\pi^{2}\alpha_{s}^{2}(t^{D_{2}})h_{f_{1}}\frac{1}{(2\pi)^{5}}\int b_{1}^{\prime}db_{1}^{\prime}\int b_{2}db_{2}$
$\int b_{3}db_{3}\int d\theta_{1}\int d\theta_{2}\exp[-S^{D_{2}}(x,x^{\prime},b,b^{\prime})]F_{1}(C,b_{2})F_{3}(A,B,D,b_{1}^{\prime},b_{3}),$ (47)
$f_{1}^{D_{3}}(q^{2}=0)=$ $\mathcal{C}M_{\Lambda_{b}}^{3}\frac{f_{\Lambda_{b}}}{8N_{c}}\frac{1}{8\sqrt{2}N_{c}}\int[dx]\int[dx^{\prime}]16\pi^{2}\alpha_{s}^{2}(t^{D_{3}})h_{f_{1}}\frac{1}{(2\pi)^{7}}\int b_{1}db_{1}\int b_{1}^{\prime}db_{1}^{\prime}\int b_{3}db_{3}$
$\int b_{3}^{\prime}db_{3}^{\prime}\int d\theta_{1}\int d\theta_{2}\int d\theta_{3}\exp[-S^{D_{3}}(x,x^{\prime},b,b^{\prime})]F_{1}\left(A,\left|\vec{b_{1}}+\vec{b_{1}^{\prime}}-\vec{b_{3}}-\vec{b_{3}^{\prime}}\right|\right)$
$F_{1}\left(B,\left|\vec{b_{3}}+\vec{b_{3}^{\prime}}\right|\right)F_{1}(C,b_{1})F_{1}(D,b_{3}^{\prime}),$ (48)
$f_{1}^{D_{4}}(q^{2}=0)=$ $\mathcal{C}M_{\Lambda_{b}}^{3}\frac{f_{\Lambda_{b}}}{8N_{c}}\frac{1}{8\sqrt{2}N_{c}}\int[dx]\int[dx^{\prime}]16\pi^{2}\alpha_{s}^{2}(t^{D_{4}})h_{f_{1}}\frac{1}{(2\pi)^{7}}\int b_{1}db_{1}\int b_{1}^{\prime}db_{1}^{\prime}\int b_{3}db_{3}$
$\int b_{3}^{\prime}db_{3}^{\prime}\int d\theta_{1}\int d\theta_{2}\int d\theta_{3}\exp[-S^{D_{4}}(x,x^{\prime},b,b^{\prime})]F_{1}\left(A,\left|\vec{b_{3}}+\vec{b_{3}^{\prime}}-\vec{b_{1}}-\vec{b_{1}^{\prime}}\right|\right)$
$F_{1}\left(B,\left|\vec{b_{3}}+\vec{b_{3}^{\prime}}\right|\right)F_{1}\left(C,\left|\vec{b_{3}}+\vec{b_{3}^{\prime}}-\vec{b_{1}}\right|\right)F_{1}(D,b_{3}),$ (49)
$f_{1}^{D_{5}}(q^{2}=0)=$ $\mathcal{C}M_{\Lambda_{b}}^{3}\frac{f_{\Lambda_{b}}}{8N_{c}}\frac{1}{8\sqrt{2}N_{c}}\int[dx]\int[dx^{\prime}]16\pi^{2}\alpha_{s}^{2}(t^{D_{5}})h_{f_{1}}\frac{1}{(2\pi)^{7}}\int b_{1}db_{1}\int b_{1}^{\prime}db_{1}^{\prime}\int b_{2}db_{2}$
$\int b_{2}^{\prime}db_{2}^{\prime}\int d\theta_{1}\int d\theta_{2}\int d\theta_{3}\exp[-S^{D_{5}}(x,x^{\prime},b,b^{\prime})]F_{1}\left(A,\left|\vec{b_{2}}+\vec{b_{2}^{\prime}}-\vec{b_{1}}-\vec{b_{1}^{\prime}}\right|\right)$
$F_{1}\left(B,\left|\vec{b_{2}}+\vec{b_{2}^{\prime}}\right|\right)F_{1}\left(C,\left|\vec{b_{2}}+\vec{b_{2}^{\prime}}-\vec{b_{1}}\right|\right)F_{1}(D,b_{2}),$ (50)
$f_{1}^{D_{6}}(q^{2}=0)=$ $\mathcal{C}M_{\Lambda_{b}}^{3}\frac{f_{\Lambda_{b}}}{8N_{c}}\frac{1}{8\sqrt{2}N_{c}}\int[dx]\int[dx^{\prime}]16\pi^{2}\alpha_{s}^{2}(t^{D_{6}})h_{f_{1}}\frac{1}{(2\pi)^{7}}\int b_{1}db_{1}\int b_{1}^{\prime}db_{1}^{\prime}\int b_{2}db_{2}$
$\int b_{2}^{\prime}db_{2}^{\prime}\int d\theta_{1}\int d\theta_{2}\int d\theta_{3}\exp[-S^{D_{6}}(x,x^{\prime},b,b^{\prime})]F_{1}\left(A,\left|\vec{b_{2}}+\vec{b_{2}^{\prime}}-\vec{b_{1}}-\vec{b_{1}^{\prime}}\right|\right)$
$F_{1}\left(B,\left|\vec{b_{2}}+\vec{b_{2}^{\prime}}\right|\right)F_{1}(C,b_{2}^{\prime})F_{1}(D,b_{1}),$ (51)
$f_{1}^{D_{7}}(q^{2}=0)=$ $\mathcal{C}M_{\Lambda_{b}}^{3}\frac{f_{\Lambda_{b}}}{8N_{c}}\frac{1}{8\sqrt{2}N_{c}}\int[dx]\int[dx^{\prime}]16\pi^{2}\alpha_{s}^{2}(t^{D_{7}})h_{f_{1}}\frac{1}{(2\pi)^{5}}\int b_{2}db_{2}\int b_{3}db_{3}$
$\int b_{3}^{\prime}db_{3}^{\prime}\int d\theta_{1}\int d\theta_{2}\exp[-S^{D_{7}}(x,x^{\prime},b,b^{\prime})]F_{1}(D,b_{3}^{\prime})F_{3}\left(A,C,B,\left|\vec{b_{3}}+\vec{b_{3}^{\prime}}\right|,b_{2}\right),$ (52)
$f_{1}^{D_{8}}(q^{2}=0)=$ $\mathcal{C}M_{\Lambda_{b}}^{3}\frac{f_{\Lambda_{b}}}{8N_{c}}\frac{1}{8\sqrt{2}N_{c}}\int[dx]\int[dx^{\prime}]16\pi^{2}\alpha_{s}^{2}(t^{D_{8}})h_{f_{1}}\frac{1}{(2\pi)^{5}}\int b_{2}db_{2}\int b_{2}^{\prime}db_{2}^{\prime}$
$\int b_{3}db_{3}\int d\theta_{1}\int d\theta_{2}\exp[-S^{D_{8}}(x,x^{\prime},b,b^{\prime})]F_{1}(C,b_{2}^{\prime})F_{3}\left(B,D,A,\left|\vec{b_{2}}+\vec{b_{2}^{\prime}}\right|,b_{3}\right),$ (53)
$f_{1}^{D_{9}}(q^{2}=0)=$ $\mathcal{C}M_{\Lambda_{b}}^{3}\frac{f_{\Lambda_{b}}}{8N_{c}}\frac{1}{8\sqrt{2}N_{c}}\int[dx]\int[dx^{\prime}]16\pi^{2}\alpha_{s}^{2}(t^{D_{9}})h_{f_{1}}\frac{1}{(2\pi)^{7}}\int b_{1}db_{1}\int b_{1}^{\prime}db_{1}^{\prime}\int b_{3}db_{3}$
$\int b_{3}^{\prime}db_{3}^{\prime}\int d\theta_{1}\int d\theta_{2}\int d\theta_{3}\exp[-S^{D_{9}}(x,x^{\prime},b,b^{\prime})]F_{1}\left(A,\left|\vec{b_{3}}+\vec{b_{3}^{\prime}}-\vec{b_{1}}-\vec{b_{1}^{\prime}}\right|\right)$
$F_{1}\left(B,\left|\vec{b_{3}}+\vec{b_{3}^{\prime}}\right|\right)F_{1}\left(C,\left|\vec{b_{3}}+\vec{b_{3}^{\prime}}-\vec{b_{1}^{\prime}}\right|\right)F_{1}(D,b_{3}^{\prime}),$ (54)
$f_{1}^{D_{10}}(q^{2}=0)=$ $\mathcal{C}M_{\Lambda_{b}}^{3}\frac{f_{\Lambda_{b}}}{8N_{c}}\frac{1}{8\sqrt{2}N_{c}}\int[dx]\int[dx^{\prime}]16\pi^{2}\alpha_{s}^{2}(t^{D_{10}})h_{f_{1}}\frac{1}{(2\pi)^{7}}\int b_{1}db_{1}\int b_{1}^{\prime}db_{1}^{\prime}\int b_{3}db_{3}$
$\int b_{3}^{\prime}db_{3}^{\prime}\int d\theta_{1}\int d\theta_{2}\int d\theta_{3}\exp[-S^{D_{10}}(x,x^{\prime},b,b^{\prime})]F_{1}\left(A,\left|\vec{b_{3}}+\vec{b_{3}^{\prime}}-\vec{b_{1}}-\vec{b_{1}^{\prime}}\right|\right)$
$F_{1}\left(B,\left|\vec{b_{3}}+\vec{b_{3}^{\prime}}\right|\right)F_{1}(C,b_{1}^{\prime})F_{1}(D,b_{3}),$ (55)
$f_{1}^{D_{11}}(q^{2}=0)=$ $\mathcal{C}M_{\Lambda_{b}}^{3}\frac{f_{\Lambda_{b}}}{8N_{c}}\frac{1}{8\sqrt{2}N_{c}}\int[dx]\int[dx^{\prime}]16\pi^{2}\alpha_{s}^{2}(t^{D_{11}})h_{f_{1}}\frac{1}{(2\pi)^{7}}\int b_{1}db_{1}\int b_{1}^{\prime}db_{1}^{\prime}\int b_{2}db_{2}$
$\int b_{2}^{\prime}db_{2}^{\prime}\int d\theta_{1}\int d\theta_{2}\int d\theta_{3}\exp[-S^{D_{11}}(x,x^{\prime},b,b^{\prime})]F_{1}\left(A,\left|\vec{b_{2}}+\vec{b_{2}^{\prime}}-\vec{b_{1}}-\vec{b_{1}^{\prime}}\right|\right)$
$F_{1}\left(B,\left|\vec{b_{2}}+\vec{b_{2}^{\prime}}\right|\right)F_{1}(C,b_{2}^{\prime})F_{1}\left(D,\left|\vec{b_{2}}+\vec{b_{2}^{\prime}}-\vec{b_{1}^{\prime}}\right|\right),$ (56)
$f_{1}^{D_{12}}(q^{2}=0)=$ $\mathcal{C}M_{\Lambda_{b}}^{3}\frac{f_{\Lambda_{b}}}{8N_{c}}\frac{1}{8\sqrt{2}N_{c}}\int[dx]\int[dx^{\prime}]16\pi^{2}\alpha_{s}^{2}(t^{D_{12}})h_{f_{1}}\frac{1}{(2\pi)^{7}}\int b_{1}db_{1}\int b_{1}^{\prime}db_{1}^{\prime}\int b_{2}db_{2}$
$\int b_{2}^{\prime}db_{2}^{\prime}\int d\theta_{1}\int d\theta_{2}\int d\theta_{3}\exp[-S^{D_{12}}(x,x^{\prime},b,b^{\prime})]F_{1}\left(A,\left|\vec{b_{2}}+\vec{b_{2}^{\prime}}-\vec{b_{1}}-\vec{b_{1}^{\prime}}\right|\right)$
$F_{1}\left(B,\left|\vec{b_{2}}+\vec{b_{2}^{\prime}}\right|\right)F_{1}(C,b_{2})F_{1}(D,b_{1}^{\prime}),$ (57)
$f_{1}^{D_{13}}(q^{2}=0)=$ $\mathcal{C}M_{\Lambda_{b}}^{3}\frac{f_{\Lambda_{b}}}{8N_{c}}\frac{1}{8\sqrt{2}N_{c}}\int[dx]\int[dx^{\prime}]16\pi^{2}\alpha_{s}^{2}(t^{D_{13}})h_{f_{1}}\frac{1}{(2\pi)^{5}}\int b_{2}db_{2}\int b_{2}^{\prime}db_{2}^{\prime}$
$\int b_{3}db_{3}\int d\theta_{1}\int d\theta_{2}\exp[-S^{D_{13}}(x,x^{\prime},b,b^{\prime})]F_{2}\left(A,B,\left|\vec{b_{2}}+\vec{b_{2}^{\prime}}\right|\right)F_{1}(C,b_{2}^{\prime})F_{1}(D,b_{3}),$ (58)
$f_{1}^{D_{14}}(q^{2}=0)=$ $\mathcal{C}M_{\Lambda_{b}}^{3}\frac{f_{\Lambda_{b}}}{8N_{c}}\frac{1}{8\sqrt{2}N_{c}}\int[dx]\int[dx^{\prime}]16\pi^{2}\alpha_{s}^{2}(t^{D_{14}})h_{f_{1}}\frac{1}{(2\pi)^{5}}\int b_{2}db_{2}\int b_{2}^{\prime}db_{2}^{\prime}$
$\int b_{3}db_{3}\int d\theta_{1}\int d\theta_{2}\exp[-S^{D_{14}}(x,x^{\prime},b,b^{\prime})]F_{2}\left(A,B,\left|\vec{b_{2}}+\vec{b_{2}^{\prime}}\right|\right)F_{1}(C,b_{2})F_{1}(D,b_{3}^{\prime}),$ (59)
where $\mathcal{C}=8/3$ is the color factor and $\equiv dx_{1}dx_{2}dx_{3}\delta(1-x_{1}-x_{2}-x_{3}), [dx][dx^{\prime}]$ is defined analogously, the auxiliary functions $A,B,C$ and $D$ in Table 8 are related to the denominators of the four propagators in each diagram. The exponent $S^{D_{i}}$ is the sum of the total exponents from the $\Lambda_{b}$ and $\Lambda$ wave functions with the hard scale $t^{D_{i}}$ involved in the diagram $D_{i}$. The functions $F_{1},F_{2}$ and $F_{3}$ are written, in terms of the Fourier integrals, as
$F_{1}(A,b)=$ $\int d^{2}k_{T}\frac{e^{i\textbf{k}_{T}\cdot\textbf{b}}}{k^{2}+A}=2\pi\left\{K_{0}(\sqrt{A}b)\theta(A)+\frac{\pi i}{2}\left[J_{0}(\sqrt{-A}b)+iN_{0}(\sqrt{-A}b)\right]\theta(-A)\right\},$ (60)
$F_{2}(A,B,b)=$ $\int d^{2}k_{T}\frac{e^{i\textbf{k}_{T}\cdot\textbf{b}}}{(k^{2}+A)(k^{2}+B)}$
$=$ $\pi\int_{0}^{1}dz\frac{b}{\sqrt{|Z_{1}|}}\left\{K_{1}(\sqrt{Z_{1}}b)\theta(Z_{1})+\frac{\pi}{2}\left[N_{1}(\sqrt{-Z_{1}}b)-iJ_{1}(\sqrt{-Z_{1}}b)\right]\theta(-Z_{1})\right\},$ (61)
$F_{3}(A,B,C,b_{1},b_{2})=$ $\int d^{2}k_{1T}\int d^{2}k_{2T}\frac{e^{i(\textbf{k}_{1T}\cdot\textbf{b}_{1}+\textbf{k}_{2T}\cdot\textbf{b}_{2})}}{(k_{1}^{2}+A)(k_{2}^{2}+B)((k_{1}+k_{2})^{2}+C)}$
$=$ $\pi^{2}\int_{0}^{1}\frac{dz_{1}dz_{2}}{z_{1}(1-z_{1})}\frac{\sqrt{X_{2}}}{\sqrt{|Z_{2}|}}$
$\times\left\{K_{1}(\sqrt{X_{2}Z_{2}})\theta(Z_{2})+\frac{\pi}{2}\left[N_{1}(\sqrt{-X_{2}Z_{2}})-iJ_{1}(\sqrt{-X_{2}Z_{2}})\right]\theta(-Z_{2})\right\},$ (62)
in which $J_{n}$ ($N_{n}$) is the Bessel function of the first (second) kind, $K_{n}$ follows the relation
$K_{n}(-iz)=\frac{\pi i}{2}e^{(in\pi)/2}\left[J_{n}(z)+iN_{n}(z)\right],$ (63)
and the variables $Z_{1}$, $Z_{2}$ and $X_{2}$ are given by
$Z_{1}=$ $Az+B(1-z),$ (64)
$Z_{2}=$ $A(1-z_{2})+\frac{z_{2}}{z_{1}(1-z_{1})}\left[B(1-z_{1})+Cz_{1}\right],$ (65)
$X_{2}=$ $(b_{1}-z_{1}b_{2})^{2}+\frac{z_{1}(1-z_{1})}{z_{2}}b_{2}^{2},$ (66)
with Feynman parameters $z$’s. Because $A$, $Z_{1}$, $Z_{2}$ and $X_{2}$ are all positive, the imaginary parts of the above functions $F_{1,2,3}$ do not contribute.
Table 8: Auxiliary functions $A,B,C$ and $D$. An overall coefficient $m_{\Lambda_{b}}^{2}$ is implicit for the entries.
A B C D
$D_{1}$ $1-x_{1}^{\prime}$ $x_{2}+x_{3}^{\prime}-x_{2}x_{3}^{\prime}$ $x_{2}x_{2}^{\prime}$ $x_{3}x_{3}^{\prime}$
$D_{2}$ $1-x_{1}^{\prime}$ $x_{2}^{\prime}+x_{3}-x_{2}^{\prime}x_{3}$ $x_{2}x_{2}^{\prime}$ $x_{3}x_{3}^{\prime}$
$D_{3}$ $1-x_{1}^{\prime}$ $\left(1-x_{1}^{\prime}\right)x_{3}$ $\left(1-x_{1}\right)\left(1-x_{1}^{\prime}\right)$ $x_{3}x_{3}^{\prime}$
$D_{4}$ $1-x_{1}^{\prime}$ $x_{3}^{\prime}\left(1-x_{1}\right)$ $\left(1-x_{1}\right)\left(1-x_{1}^{\prime}\right)$ $x_{3}x_{3}^{\prime}$
$D_{5}$ $1-x_{1}^{\prime}$ $x_{2}^{\prime}\left(1-x_{1}\right)$ $\left(1-x_{1}\right)\left(1-x_{1}^{\prime}\right)$ $x_{2}x_{2}^{\prime}$
$D_{6}$ $1-x_{1}^{\prime}$ $x_{2}\left(1-x_{1}^{\prime}\right)$ $x_{2}x_{2}^{\prime}$ $\left(1-x_{1}\right)\left(1-x_{1}^{\prime}\right)$
$D_{7}$ $x_{3}\left(1-x_{2}^{\prime}\right)$ $1-x_{1}$ $x_{2}x_{2}^{\prime}$ $x_{3}x_{3}^{\prime}$
$D_{8}$ $x_{2}\left(1-x_{3}^{\prime}\right)$ $1-x_{1}$ $x_{2}x_{2}^{\prime}$ $x_{3}x_{3}^{\prime}$
$D_{9}$ $1-x_{1}$ $x_{3}\left(1-x_{1}^{\prime}\right)$ $\left(1-x_{1}\right)\left(1-x_{1}^{\prime}\right)$ $x_{3}x_{3}^{\prime}$
$D_{10}$ $1-x_{1}$ $x_{3}^{\prime}\left(1-x_{1}\right)$ $\left(1-x_{1}\right)\left(1-x_{1}^{\prime}\right)$ $x_{3}x_{3}^{\prime}$
$D_{11}$ $1-x_{1}$ $x_{2}\left(1-x_{1}^{\prime}\right)$ $x_{2}x_{2}^{\prime}$ $\left(1-x_{1}\right)\left(1-x_{1}^{\prime}\right)$
$D_{12}$ $1-x_{1}$ $x_{2}^{\prime}\left(1-x_{1}\right)$ $x_{2}x_{2}^{\prime}$ $\left(1-x_{1}\right)\left(1-x_{1}^{\prime}\right)$
$D_{13}$ $x_{2}\left(1-x_{3}^{\prime}\right)$ $x_{2}+x_{3}^{\prime}-x_{2}x_{3}^{\prime}$ $x_{2}x_{2}^{\prime}$ $x_{3}x_{3}^{\prime}$
$D_{14}$ $x_{3}\left(1-x_{2}^{\prime}\right)$ $x_{2}^{\prime}+x_{3}-x_{2}^{\prime}x_{3}$ $x_{2}x_{2}^{\prime}$ $x_{3}x_{3}^{\prime}$